OpenStudy (anonymous):

Suppose two sides of a square are given as x + 3 and 12 - x. What is its area?

OpenStudy (anonymous):

do you need an explanation or just the answer

OpenStudy (anonymous):

You know that the sides are equal - since it is a square

OpenStudy (anonymous):

both step by step please

OpenStudy (anonymous):

Area=(base)(height)

OpenStudy (anonymous):

ok hold on please

OpenStudy (anonymous):

x+3 = 12-x

OpenStudy (anonymous):

Soo x equals 4.5

OpenStudy (anonymous):

4.5+3=12-4.5 7.5*7.5 56.25

OpenStudy (anonymous):

Aaaand now YOU COMPUTE the sides themselves from the assumed lengths .....

OpenStudy (anonymous):

What is the gratitude here ;]